A plane took 1 hour longer to travel 560 miles on the first portion of a flight than it took to fly 480 miles on the second portion. If the speed was the same for each portion, what was the flying time for the second part of the trip?
My answer is 7 hours am I right?

Answer by checkley75(3666)

X/480=(X+1)/560
560X=480X+480
560X-480X=480
80X=480
X=480/80
X=6 ANSWER FOR THE 480 MILE FLIGHT.
6+1=7 HOURS FOR THE 560 MILE FLIGHT.